Could someone help me out?
How much of a 40% antifreeze solution must a mechanical mix with an 80% antifreeze solution if 20 L of a 50% antifreeze solution are need?
I tried the following and was marked wrong.
Let c represent the 40% solution
___ y _____________ 80% ________
x + y= 20
40x+80y= 50
------------
40x+40y=8
40x+40y=.5
-4y=7.5
y=18.75
------------
x+18.75=20
------------
x=1.25 y= 18.75

Answer by stanbon(75887)

How much of a 40% antifreeze solution must a mechanic mix with an 80% antifreeze solution if 20 L of a 50% antifreeze solution are needed?
------------------------
active + active = active
0.40x + 0.80(20-x) = 0.50*20
40x + 1600 - 80x = 1000
-40x = -600
x = 15 liters (amount of 40% antifreeze needed in the mixture)
-------------
20-x = 5 liters (amount of 80% antifreeze needed in the mixture)
=========================
Cheers,
Stan H.